Biscuit is a Bakery. Until now Biscuit has been an all-equity firm, with asset fully depreciated, and its current expected cash flow to equity holders is 1.5 million NOK. The firm now undertakes a perpetual debt for 300,000 NOK to buy a new oven that will be depreciated straight-line over three years. The expected return on unlevered equity is 10.05% and the cost of debt is 3.00%. The tax rate on corporate earnings is 26%. The expected EBITDA of the company is not affected by the purchase of the new oven and is constant over time. The depreciation tax shield is as risky as the company’s assets.

What is the value of Biscuit's assets?

A. 14,466,493
B. 14,989,854
C. 15,067,974
D. 15,125,690
The correct answer is C
Answer #1
PV of Expected FFE = 1500000/10.05% = $          1,49,25,373
PV of depreciation tax shield for 3 years = 100000*26%*(1.1005^3-1)/(0.1005*1.1005^3) = $                    64,601
Value of the unlevered firm $          1,49,89,974
Value tax shield on interest = 300000*26% = $                    78,000
Value of the levered firm (assets) $          1,50,67,974
Answer: Option C $15,067,974